I have to solve this system.
-2(x+y)-3=5x+3y-10
-3y-(x+2)=3y-5x+5
I have worked it several times and keep getting different answers.
Please help!

Answer by venugopalramana(3286)

I have to solve this system.
-2(x+y)-3=5x+3y-10
-2X-2Y-3=5X+3Y-10
7X+5Y=7...............................I
-3y-(x+2)=3y-5x+5
-3Y-X-2=3Y-5X+5
4X-6Y=7..................................II
6*EQN.I+5*EQN.II
42X+30Y+20X-30Y=42+35=77
62X=77
X=77/62
4*77/62-6Y=7
(308-434)/62=6Y
Y=-126/372=-63/186=-21/62
I have worked it several times and keep getting different answers.
Please help!